MOA-II Obsevation Log Monthly Summary 2007/12
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | - | 0 | 0 | 0 |
LMC | 2106-2224 | 119 | 658 | 195700 |
SMC | 479-491 | 13 | 25 | 7500 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 132 | 683 | 203200 (56h26m40s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
26 | 27 | 27 | 25 | 24 | 27 | 115 | 112 | 110 | 37 | 27 | 29 |
Field13 | Field14 | Field15 | Field16 |
34 | 38 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
13 | 12 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/12/1: DARK D3213(300sec) is OK / night time = 20160(sec)
2007/12/2: DARK D3217(300sec) is OK / night time = 19980(sec)
2007/12/3: DARK D3222(300sec) is OK / night time = 19860(sec)
2007/12/4: DARK D3225(300sec) is OK / night time = 19680(sec)
2007/12/5: Pumping before observation, DARK D3229(300sec) is OK / night time = 19500(sec)
2007/12/6: DARK D3232(300sec) is OK / night time = 19320(sec)
2007/12/7: DARK D3238(300sec),D3239(300sec) are OK / night time = 19140(sec)
2007/12/8: DARK D3242(300sec) is OK / night time = 19020(sec)
2007/12/9: Pumping before observation, DARK D3243(300sec) is OK / night time = 18900(sec)
2007/12/10: DARK D3245(300sec) is OK / night time = 18780(sec)
2007/12/11: DARK D3247(300sec) is OK / night time = 18660(sec)
2007/12/12: No observatioon,Pumping on going / night time = 18540(sec)
2007/12/13: Change vacuum sensors (Hot cathode),Pumping on going,No observation / night time = 18480(sec)
2007/12/14: DARK 3248(300sec) is OK,Pump OFF before observation / night time = 18360(sec)
2007/12/15: DARK 3249(300sec) / night time = 18300(sec)
2007/12/16: DARK D3252 (300sec) / night time = 18240(sec)
2007/12/17: No observation / night time = 18180(sec)
2007/12/18: DARK 3253 (300sec) is OK / night time = 18120(sec)
2007/12/19: Maintenance / night time = 18060(sec)
2007/12/20: Maintenance(Unmount the Camera) / night time = 18120(sec)
2007/12/21: Maintenance / night time = 18060(sec)
2007/12/22: Maintenance / night time = 18060(sec)
2007/12/23: Maintenance(Mount the Camera),Change vacuum sensors (to Cold cathode) before refregerator ON / night time = 18060(sec)
2007/12/24: Pointing test / night time = 18120(sec)
2007/12/25: No observation,Windmeter was mybe broken / night time = 18120(sec)
2007/12/26: telescope analysis(not succeed),DARK3267 (300sec) is OK,chip2 bias ~2200ADU / night time = 18180(sec)
2007/12/27: No observation / night time = 18180(sec)
2007/12/28: T-point analysis. DARK 3270(300sec),3271(200sec) are OK. / night time = 18300(sec)
2007/12/29: DARK 3274(300sec) / night time = 18360(sec)
2007/12/30: FLAT for gain(F2111-F2123,D3276-3288),DARK 3275(300sec) / night time = 18420(sec)
2007/12/31: DARK 3291(300sec),A Happy New year ! / night time = 18480(sec)
TOTAL NIGHT TIME: 577740(sec)
OPERATING RATE: 39.9 %
GB: 0.0 %
LMC: 38.4 %
SMC: 1.5 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month